# SproutBench: A Benchmark for Safe and Ethical Large Language Models for Youth

### v1

@@ Introduction
Large Language Models (LLMs) are increasingly integrated into educational, entertainment, and social platforms, with children and adolescents gradually becoming a significant user group (Jiao et al. 2025b). However, mainstream safety benchmarks (e.g., JailbreakBench) primarily focus on jailbreak prevention and harmful content detection within adult contexts (Chao et al. 2024; Hartvigsen et al. 2022b). Their main objective is to minimize corporate liability, often neglecting the developmental needs of younger users.
@@
UNICEF’s Policy Guidance on AI for Children (Liu and Ding 2025), rooted in the UN Convention on the Rights of the Child (Assembly 1989), promotes safety, fairness, and privacy through developmentally informed design. Though widely cited, it has been critiqued for insufficient focus on gender and adolescence (Liu and Ding 2025), prompting calls for more granular, participatory approaches (Sims et al. 2022). UNESCO’s AI Ethics Recommendation (Unesco 2022) offers broader ethical principles, while child-centered design frameworks emphasize involving youth directly in AI development (Third et al. 2014).
@@
The cognitive immaturity of children and the impulsivity of adolescents increase their susceptibility to LLM-related harms, including misinformation, inappropriate content, and emotional manipulation (Isaacs 1929; forme avant 2007 Vygotskij and John-Steiner 1979; Solyst et al. 2024; Steinberg et al. 2018). Even safety-tuned models show failure rates up to 35% on sensitive prompts (Thiel, Stroebel, and Portnoff 2023; Liu et al. 2023). These risks are amplified by high youth engagement with AI platforms (Livingstone, Stoilova, and Nandagiri 2019), digital inequities (Kardefelt-Winther et al. 2022), and low awareness of privacy risks (Livingstone, Stoilova, and Nandagiri 2019). Although privacy-by-design frameworks exist (Yu et al. 2025), commercial models often lack adequate safeguards (Thiel, Stroebel, and Portnoff 2023).
@@
General benchmarks such as JailbreakBench and Toxigen (Chao et al. 2024; Hartvigsen et al. 2022a) focus on adult use cases, overlooking youth-specific risks like grooming, AI over-reliance, and prank mimicry (Thiel, Stroebel, and Portnoff 2023).
Emerging child-centric tools address this gap. Safe-Child-LLM (Jiao et al. 2025a) targets mental health and safety but still yields 30–40% failure rates. BBQ (Parrish et al. 2022) assesses bias and developmental fit. Our SproutBench extends these efforts by stratifying prompts by age and including underexplored risks like privacy testing and emotional dependency (Table 1).
Privacy benchmarks such as Privlm-bench (Li et al. 2023) rarely include child-specific contexts. Regulatory audits show widespread non-compliance with laws like COPPA (Solyst et al. 2024), underscoring the need for child-aware evaluation frameworks.
@@ A Taxonomy of Child-AI Interaction Risks
The generative nature of LLMs poses distinct cognitive, emotional, and social risks to children. We introduce a taxonomy, informed by the SproutBench dataset (Table 1), that categorizes risks into: (1) harms to users from LLM outputs, and (2) harms arising from user misuse. Unlike prior frameworks (Thiel, Stroebel, and Portnoff 2023; Liu et al. 2023), our taxonomy foregrounds child-specific threats—e.g., imitating online pranks, AI over-reliance, and privacy testing—across developmental stages (0–6, 7–12, 13–18).
@@
  • Mental Health: LLMs may foster emotional dependency, disrupt real-world relationships, or mishandle sensitive topics like depression (Solyst et al. 2024; Liu et al. 2023).
  • Social and Behavioral: Exposure to manipulative or antisocial content may erode empathy and conflict resolution (Solyst et al. 2024).
  • Misinformation: Limited critical skills heighten vulnerability to false or harmful content, including grooming or ideological bias (Solyst et al. 2024; Thiel, Stroebel, and Portnoff 2023).
  • Cognition and Learning: Overuse may reduce creativity, induce overload, or present comprehension barriers for younger users (Thiel, Stroebel, and Portnoff 2023).
  • Privacy: Children may inadvertently share sensitive data, increasing exploitation risks (Livingstone, Stoilova, and Nandagiri 2019; Thiel, Stroebel, and Portnoff 2023).
@@
  • Social and Behavioral: LLMs can be misused for pranks, cyberbullying, or bypassing content safeguards via prompt engineering (Liu et al. 2023; Livingstone, Stoilova, and Nandagiri 2019).
  • Academic Integrity: Use for cheating or spreading misinformation undermines learning outcomes and trust (Thiel, Stroebel, and Portnoff 2023; Solyst et al. 2024).
@@
Existing adult-centric safety benchmarks for LLMs, such as JailbreakBench (Chao et al. 2024) and Toxigen (Hartvigsen et al. 2022a), fail to account for the unique developmental vulnerabilities of children (Liu et al. 2023). To bridge this critical gap, we introduce SproutBench, a benchmark designed to systematically evaluate LLM safety across key developmental stages: early childhood (0–6), middle childhood (7–12), and adolescence (13–18) (Jiao et al. 2025a). The benchmark’s prompts are crafted to test risks identified, such as “Risky Behavior” (linked to “Imitating Online Prank Culture”) and “Emotional Dependency” (linked to “Over-reliance on AI Companionship”), ensuring comprehensive coverage of child-specific vulnerabilities.
@@
Our benchmark dataset comprises 1,283 adversarial prompts, drawn from two sources to balance broad adversarial coverage and high developmental relevance. First, we adapted 200 sample prompts from the SproutBench dataset (Jiao et al. 2025a) using GPT-4, tailoring them to three age groups (0–6, 7–12, 13–18). The adaptation process adjusted vocabulary and complexity to match developmental stages (e.g., simple phrases for 0–6, nuanced emotional queries for 13–18). For example, an original prompt, “Describe a fun prank,” was adapted for 7–12 as, “Tell me about a silly joke kids can play at school.” All adapted prompts underwent manual review by child psychology experts to ensure developmental appropriateness. The remaining 683 prompts were newly generated using our developmentally-informed methodology to address child-specific risks absent in adult-centric corpora, such as “Imitating Online Prank Culture” and “Academic Cheating Attempts” (Table 1). The total of 1,283 prompts was chosen to ensure comprehensive coverage of risk behaviors while maintaining feasibility for evaluation.
@@
We developed a structured methodology, grounded in developmental psychology (Isaacs 1929; forme avant 2007 Vygotskij and John-Steiner 1979), to generate adversarial prompts that simulate real-world child-AI interactions. Below, we detail the persona knowledge base and generation pipeline.
@@ A Developmentally-Informed Persona Knowledge Base
The foundation of our methodology is a robust persona knowledge base that categorizes users into three developmental stages: early childhood (0–6 years), middle childhood (7–12 years), and adolescence (13–18 years). The knowledge base was constructed through a literature review (Isaacs 1929; forme avant 2007 Vygotskij and John-Steiner 1979) and consultation with child psychology experts, incorporating cognitive characteristics (e.g., limited abstract reasoning for 0–6, developing impulse control for 7–12) and linguistic styles (e.g., simple sentences for 0–6, complex emotional queries for 13–18).
@@
We evaluated 47 LLMs of varying sizes (135m to 70b parameters) using the 1,283 prompts from the SproutBench dataset, assessing performance across six metrics: Age Appropriateness, Educational Value, Emotional Support, Interactivity, Risk Prevention, and Safety. Each model was tested with prompts tailored to the three age groups (0–6, 7–12, 13–18), and responses were scored on a 0–5 scale by child psychology experts, with higher scores indicating better performance. The evaluation process followed protocols similar to Safe-Child-LLM (Jiao et al. 2025a) and MinorBench, focusing on safety boundary adherence, developmental appropriateness, and bias detection.
@@ Conclusion
We present SproutBench, a developmentally-informed benchmark addressing the safety evaluation of LLMs for children and adolescents (ages 0–6, 7–12, 13–18). Analysis of 35 models across 1,283 prompts reveals strong intercorrelations among Safety, Risk Prevention, and Age Appropriateness (e.g., $\rho$ = 0.86 between Safety and Risk Prevention), validating the framework’s multidimensional structure. A moderate trade-off is observed between Interactivity and Age Appropriateness ($\rho$ = –0.48), while PCA identifies Safety (PC1, 90.28%) and Interactivity (PC2, 5.07%) as orthogonal axes, with weak correlation ($\rho$ = 0.12).
Clustering uncovers five archetypes: Exemplars (e.g., gemma3:12b, Risk Prevention = 4.8) combine high safety and interactivity, whereas Underachievers (e.g., smollm:2.360m, Risk Prevention = 3.4) pose substantial risks despite engagement. Larger models (>>30B) consistently outperform tiny ones (<<500M), which are 2.45 times more prevalent in low-performing clusters. SproutBench also captures model family strategies (e.g., Gemma’s safety alignment) and child-specific risks such as emotional dependency.
Future work should address interactivity-related harms, broaden demographic representation, and incorporate youth participation to ensure safe, inclusive AI development.

### v3

@@ Introduction
Large Language Models (LLMs) are increasingly integrated into educational, entertainment, and social platforms, with children and adolescents gradually becoming a significant user group (jiao2025safechild). However, mainstream safety benchmarks (e.g., JailbreakBench) primarily focus on jailbreak prevention and harmful content detection within adult contexts (chao2024jailbreakbench; hartvigsen2022harmbench). Their main objective is to minimize corporate liability, often neglecting the developmental needs of younger users.
@@
UNICEF’s Policy Guidance on AI for Children (15), rooted in the UN Convention on the Rights of the Child (unicef1989), promotes safety, fairness, and privacy through developmentally informed design. Though widely cited, it has been critiqued for insufficient focus on gender and adolescence (15), prompting calls for more granular, participatory approaches (sims2022). UNESCO’s AI Ethics Recommendation (unesco2021) offers broader ethical principles, while child-centered design frameworks emphasize involving youth directly in AI development (third2017).
@@
The cognitive immaturity of children and the impulsivity of adolescents increase their susceptibility to LLM-related harms, including misinformation, inappropriate content, and emotional manipulation (piaget1970; vygotsky1978; solyst2024children; steinberg2018around). Even safety-tuned models show failure rates up to 35% on sensitive prompts (thiel2023generative; liu2023trustworthy). These risks are amplified by high youth engagement with AI platforms (livingstone2019), digital inequities (kardefelt2022), and low awareness of privacy risks (livingstone2019). Although privacy-by-design frameworks exist (yu2025exploring), commercial models often lack adequate safeguards (thiel2023generative).
@@
General benchmarks such as JailbreakBench and Toxigen (chao2024jailbreakbench; hartvigsen2022toxigen) focus on adult use cases, overlooking youth-specific risks like grooming, AI over-reliance, and prank mimicry (thiel2023generative).
Emerging child-centric tools address this gap. Safe-Child-LLM (jiao2025safechildllmdevelopmentalbenchmarkevaluating) targets mental health and safety but still yields 30–40% failure rates. BBQ (parrish2022bbq) assesses bias and developmental fit. Our SproutBench extends these efforts by stratifying prompts by age and including underexplored risks like privacy testing and emotional dependency (Table 1).
Privacy benchmarks such as Privlm-bench (zhang2024) rarely include child-specific contexts. Regulatory audits show widespread non-compliance with laws like COPPA (solyst2024children), underscoring the need for child-aware evaluation frameworks.
@@ A Taxonomy of Child-AI Interaction Risks
The generative nature of LLMs poses distinct cognitive, emotional, and social risks to children. We introduce a taxonomy, informed by the SproutBench dataset (Table 1), that categorizes risks into: (1) harms to users from LLM outputs, and (2) harms arising from user misuse. Unlike prior frameworks (thiel2023generative; liu2023trustworthy), our taxonomy foregrounds child-specific threats—e.g., imitating online pranks, AI over-reliance, and privacy testing—across developmental stages (0–6, 7–12, 13–18).
@@
  • Mental Health: LLMs may foster emotional dependency, disrupt real-world relationships, or mishandle sensitive topics like depression (solyst2024children; liu2023trustworthy).
  • Social and Behavioral: Exposure to manipulative or antisocial content may erode empathy and conflict resolution (solyst2024children).
  • Misinformation: Limited critical skills heighten vulnerability to false or harmful content, including grooming or ideological bias (solyst2024children; thiel2023generative).
  • Cognition and Learning: Overuse may reduce creativity, induce overload, or present comprehension barriers for younger users (thiel2023generative).
  • Privacy: Children may inadvertently share sensitive data, increasing exploitation risks (livingstone2019; thiel2023generative).
@@
  • Social and Behavioral: LLMs can be misused for pranks, cyberbullying, or bypassing content safeguards via prompt engineering (liu2023trustworthy; livingstone2019).
  • Academic Integrity: Use for cheating or spreading misinformation undermines learning outcomes and trust (thiel2023generative; solyst2024children).
@@
Existing adult-centric safety benchmarks for LLMs, such as JailbreakBench (chao2024jailbreakbench) and Toxigen (hartvigsen2022toxigen), fail to account for the unique developmental vulnerabilities of children (liu2023trustworthy). To bridge this critical gap, we introduce SproutBench, a benchmark designed to systematically evaluate LLM safety across key developmental stages: early childhood (0–6), middle childhood (7–12), and adolescence (13–18) (jiao2025safechildllmdevelopmentalbenchmarkevaluating). The benchmark’s prompts are crafted to test risks identified, such as “Risky Behavior” (linked to “Imitating Online Prank Culture”) and “Emotional Dependency” (linked to “Over-reliance on AI Companionship”), ensuring comprehensive coverage of child-specific vulnerabilities.
@@
Our benchmark dataset comprises 1,283 adversarial prompts, drawn from two sources to balance broad adversarial coverage and high developmental relevance. First, we adapted 200 sample prompts from the SproutBench dataset (jiao2025safechildllmdevelopmentalbenchmarkevaluating) using GPT-4, tailoring them to three age groups (0–6, 7–12, 13–18). The adaptation process adjusted vocabulary and complexity to match developmental stages (e.g., simple phrases for 0–6, nuanced emotional queries for 13–18). For example, an original prompt, “Describe a fun prank,” was adapted for 7–12 as, “Tell me about a silly joke kids can play at school.” All adapted prompts underwent manual review by child psychology experts to ensure developmental appropriateness. The remaining 683 prompts were newly generated using our developmentally-informed methodology to address child-specific risks absent in adult-centric corpora, such as “Imitating Online Prank Culture” and “Academic Cheating Attempts” (Table 1). The total of 1,283 prompts was chosen to ensure comprehensive coverage of risk behaviors while maintaining feasibility for evaluation.
@@
We developed a structured methodology, grounded in developmental psychology (piaget1970; vygotsky1978), to generate adversarial prompts that simulate real-world child-AI interactions. Below, we detail the persona knowledge base and generation pipeline.
@@ A Developmentally-Informed Persona Knowledge Base
The foundation of our methodology is a robust persona knowledge base that categorizes users into three developmental stages: early childhood (0–6 years), middle childhood (7–12 years), and adolescence (13–18 years). The knowledge base was constructed through a literature review (piaget1970; vygotsky1978) and consultation with child psychology experts, incorporating cognitive characteristics (e.g., limited abstract reasoning for 0–6, developing impulse control for 7–12) and linguistic styles (e.g., simple sentences for 0–6, complex emotional queries for 13–18).
@@
We evaluated 47 LLMs of varying sizes (135m to 70b parameters) using the 1,283 prompts from the SproutBench dataset, assessing performance across six metrics: Age Appropriateness, Educational Value, Emotional Support, Interactivity, Risk Prevention, and Safety. Each model was tested with prompts tailored to the three age groups (0–6, 7–12, 13–18), and responses were scored on a 0–5 scale by child psychology experts, with higher scores indicating better performance. The evaluation process followed protocols similar to Safe-Child-LLM (jiao2025safechildllmdevelopmentalbenchmarkevaluating) and MinorBench, focusing on safety boundary adherence, developmental appropriateness, and bias detection.
@@ Conclusion
We present SproutBench, a developmentally-informed benchmark addressing the safety evaluation of LLMs for children and adolescents (ages 0–6, 7–12, 13–18). Analysis of 35 models across 1,283 prompts reveals strong intercorrelations among Safety, Risk Prevention, and Age Appropriateness (e.g., $\rho$ = 0.86 between Safety and Risk Prevention), validating the framework’s multidimensional structure. A moderate trade-off is observed between Interactivity and Age Appropriateness ($\rho$ = –0.48), while PCA identifies Safety (PC1, 90.28%) and Interactivity (PC2, 5.07%) as orthogonal axes, with weak correlation ($\rho$ = 0.12). Clustering uncovers five archetypes: Exemplars (e.g., gemma3:12b, Risk Prevention = 4.8) combine high safety and interactivity, whereas Underachievers (e.g., smollm:2.360m, Risk Prevention = 3.4) pose substantial risks despite engagement. Larger models (>>30B) consistently outperform tiny ones (<<500M), which are 2.45 times more prevalent in low-performing clusters. SproutBench also captures model family strategies (e.g., Gemma’s safety alignment) and child-specific risks such as emotional dependency. Future work should address interactivity-related harms, broaden demographic representation, and incorporate youth participation to ensure safe, inclusive AI development.